jack is twice as old as jill and six years younger than kathy. if jill is h years old, how old is cathy in terms of h?
Found 2 solutions by josgarithmetic, ankor@dixie-net.com:
Answer by josgarithmetic(39620)

Following directly from the description:
Jack         2h
Jill          h
Kathy        2h+6



-
Who is "cathy"?

Answer by ankor@dixie-net.com(22740)

jack is twice as old as jill and six years younger than kathy.
if jill is h years old, how old is cathy in terms of h?
"
let j =Jack's age, let h = Jill's age, let k - Kathy's age
write an equation for each statement
" jack is twice as old as jill"
j = 2h
" and six years younger than kathy."
j = k-6
:
if jill is h years old, how old is kathy in terms of h?
replace j with 2h in the above equation
2h = k-6
2h+6 = k
or
k = 2h+6